Workshop on Representation Theory Bielefeld, June 28–30, 1999 LIST OF PARTICIPANTS

1. Bernd ACKERMANN (Stuttgart), ackermbd@cip.mathematik.uni-stuttgart.de
2. Manfred BAUCH (Bayreuth), Manfred.Bauch@uni-bayreuth.de
3. Thomas BRÜSTLE (Bielefeld), bruestle@ mathematik.uni-bielefeld.de
4. Gabriella D’ESTE (Milano), deste@vmimat.mat.unimi.it
5. Richard DIPPER (Stuttgart), rdipper@mathematik.uni-stuttgart.de
6. Peter DRÄXLER (Bielefeld), draexler@mathematik.uni-bielefeld.de
7. Rolf FARNSTEINER (Bielefeld), rolf@mathematik.uni-bielefeld.de
8. Jörg FELDVOSS (Bielefeld), feldvoss@mathematik.uni-bielefeld.de
9. Jerome GERMONI (Lyon), germoni@desargues.univ-lyon1.fr
10. Jin Yun GUO (Bielefeld/Changsha), gjy@sparc2.hunnu.edu.cn
11. Anne HENKE (Oxford), henke@maths.ox.ac.uk
12. Lutz HILLE (Hamburg), hille@math.uni-hamburg.de
13. Jiuzhao HUA (Bielefeld), sfb11@mathematik.uni-bielefeld.de
14. Gordon JAMES (London), g.james@ic.ac.uk
15. Steffen KÖNIG (Bielefeld), koenig@mathematik.uni-bielefeld.de
16. Shigeo KOSHITANI (Essen/Chiba), koshitan@math.7.chiba-u.ac.jp
17. Amrey KRAUSE (Paderborn), akrause@math.uni-paderborn.de
18. Mathias KÜNZER (Bielefeld), kuenzer@mathematik.uni-bielefeld.de
19. Bernard LECLERC (Caen), leclerc@math.unicaen.fr
20. Nikolaos MARMARIDIS (Ioannina), nmarmar@cc.ioi.gr
21. Andrew MATHAS (Sydney), mathas@maths.usyd.edu.au
22. Volodymyr MAZORCHUK (Kiew/Bielefeld), mazor@mathematik.uni-bielefeld.de
23. Rainer NÖRENBERG (Bielefeld), noerenbg@mathematik.uni-bielefeld.de
24. Alexander POLISHCHUK (Harvard), apolish@yahoo.com
25. Rudolf RENTSCHLER (Paris), rent@ccr.jussieu.fr
26. Claus Michael RINGEL (Bielefeld), ringel@mathematik.uni-bielefeld.de
27. Gerhard RÖHRLE (Bielefeld), roehrle@mathematik.uni-bielefeld.de
28. Bert SEVENHANT (Diepenbeek/Hasselt), bert.sevenhant@luc.ac.be
29. Andreij SKOWRONSKI (Torun), skowron@math.uni.torun.pl
30. Inger Heidi SLUNGÅRD (Trondheim/Bielefeld), ingersl@math.ntnu.no

31. Czaba SZANTO (Cluji/Bielefeld), szanto@mathematik.uni-bielefeld.de
32. Marcus SCHLIEMANN (Stuttgart), mschliemann@mathematik.uni-stuttgart.de
33. Youjun TAN (Bielefeld), youtan@mathematik.uni-bielefeld.de
34. Michela VARAGNOLO (Cergy Pontoise), varagnolo@math.pst.u-cergy.fr
35. Eric VASSEROT (Paris), vasserot@u-cergy.fr
36. Jie XIAO (Beijing/Bielefeld), xiaojie@bnu.edu.cn
37. Shilin YANG (Bielefeld), slyang@mathematik.uni-bielefeld.de
38. Jiachen YE (Bielefeld/Shanghai), jcyed@ online.sh.cn